# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ A.4 Literature review
    The training dataset $\mathcal{T}=\{(\bm{x}^{(i)},\hat{\bm{o}}^{(i)}_{\bm{a}^{*}})\}_{i=1}^{n}$ with the size $n$ is collected based on $f_{\bm{a}^{*}}$, i.e., for the $i$-th training example, the classical input $\bm{x}^{(i)}$ randomly sampled from $[-\pi,\pi]^{d}$ is fed into the quantum system described by $f_{\bm{a}^{*}}$ and the statistical estimation of $f_{\bm{a}^{*}}(\bm{x}^{(i)},O)$ is obtained by $T$ shots with $\hat{\bm{o}}^{(i)}_{\bm{a}^{*}}=\frac{1}{T}\sum_{t}\hat{\bm{o}}^{(i)}_{\bm{a}^{*},t}$ and $\mathbb{E}(\hat{\bm{o}}^{(i)}_{\bm{a}^{*}})=f_{\bm{a}^{*}}(\bm{x}^{(i)},O)$;
@@
Remark. For ease of analysis, in the proof of Theorem B.1, we consider the shot number $T$ is sufficiently large such that the measured results can be approximated by the normal distribution with the mean $\mu_{\bm{a}}=f_{\bm{a}}(\bm{x})$. Besides, for each $\bm{a}$, the corresponding variance of measured results is assumed to be equal with the varied $\bm{x}$, i.e., for all $\bm{x}$, the variance is $\nu_{\bm{a}}=\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\nu_{\bm{a}}(\bm{x})$.
@@
For the generic variance $\nu^{2}$, it is derived as follows. When the state $\rho_{\bm{a}}(\bm{x})$ is measured by $O=BX$, the probability of measuring the state associated with the eigenvalue $+1$ and the state associated with the eigenvalue $-1$ is
@@
Following the results [109, Lemma 9], the right-hand side term is upper bounded by the number of training examples and the qubit count, i.e.,
@@
Consider a learner quires an $N$-qubit quantum circuits and collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[\pi,\pi]^{d}$. Suppose with high probability, there exists a learned model $h(\cdot,\cdot)$ that can achieve
@@
=$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}(\rho(\bm{x})O)\Big{|}\bm{x}\in[-\pi,\pi]^{d},\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})\right\},$ (65)
where $\bm{x}\in[-\pi,\pi]^{d}$ is the classical input control, $\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})$ refers to the set of circuit layouts formed by $d$ $\operatorname{RZ}$ gates and $G-d$ Clifford gates, and $\rho(\bm{x})$ is the quantum state generated by the bounded gates with $\bm{x}$, i.e., $\rho(\bm{x})=U(\bm{x})\rho_{0}U(\bm{x})^{\dagger}$ and the layout of $U(\bm{x})$ follows an element of $\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})$. The diversity of circuit layouts stems from the constraint that the learner can solely utilize the classical input and the corresponding measurement results to infer the target concept, where prior information regarding the quantum circuit indicates that it comprises $G$ gates, with $d$ $\operatorname{RZ}$ gates and $d-G$ CI gates. From the perspective of the learner, the quantum computer can yield many circuit layouts, leading that the function space $\mathcal{F}$ contains in total $|\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})|$ circuit layouts.
@@
[…]
Remark. Note that although the original proof of Proposition D.1 only concerns $\|O\|\leq 1$, the pertinent proof can be readily extended to the setting of $\sum_{i}\|O_{i}\|_{\infty}\leq B$ concerned in Theorem 1, and therefore we omit it here. Moreover, it only focuses on the case of $T=1$, but the results still hold for the case of $T>1$. This is because in the extreme case with $T\rightarrow\infty$, the required training data size is reduced to $n\geq{38B^{2}\log(2\mathcal{M}(\mathcal{F},4\epsilon,\mathfrak{d})/\delta)}/{\epsilon}$ [69, Equation (C66)]. In this regard, the setting of $T\geq 1$ only trivially influences the sample complexity bound (at most logarithmically). For this reason, we omit the relevant analysis in our proof.
@@
We now derive the upper bound of $|\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})|$. The total number of possible layouts for an $N$-qubit circuit consisting with $d$ $\operatorname{RZ}$ gates and $G-d$ CI gates (i.e., $\text{CI}=\{H,S,\operatorname{CNOT}\}$) is
@@
[…]
@@
The complexity class $\mathsf{BQP}$. $\mathsf{BQP}$ (short for bounded-error quantum polynomial time) is the class of promise problems that can be solved by polynomial-time quantum computations that may have some small probability of making an error [115]. Let $A=(A_{\text{yes}},A_{\text{no}})$ be a promise problem and let $a,b:\mathbb{N}\rightarrow[0,1]$ be functions. Then $A\in\mathsf{BQP}(a,b)$ if and only if there exists a polynomial-time generated family of quantum circuits $\mathcal{Q}=\{\mathcal{Q}_{N}:N\in\mathbb{N}\}$, where each circuit $\mathcal{Q}_{N}$ takes $N$ input qubits and produces one output qubit, that satisfies the following properties: (1) if $\bm{x}\in A_{\text{yes}}$ then $\Pr(\mathcal{Q}\leavevmode\nobreak\ \text{accepts}\leavevmode\nobreak\ \bm{x})\geq a(|\bm{x}|)$, and (2) if $\bm{x}\in A_{\text{no}}$ then $\Pr(\mathcal{Q}\leavevmode\nobreak\ \text{accepts}\leavevmode\nobreak\ \bm{x})\leq b(|\bm{x}|)$. The class is defined as $\mathsf{BQP}=\mathsf{BQP}(2/3,1/3)$.
The complexity class $\mathsf{HeurP}^{\mathbb{D}}/\mathsf{poly}$. The complexity class $\mathsf{HeurP}^{\mathbb{D}}/\mathsf{poly}$ refers to the class of languages $\mathsf{L}$ that can be decided in $\mathsf{HeurP}/\mathsf{poly}$ under the distribution $\mathbb{D}$, i.e. $\mathsf{HeurP}^{\mathbb{D}}/\mathsf{poly}=\{\mathsf{L}|(\mathsf{L},\mathbb{D})\in\mathsf{HeurP}/\mathsf{poly}\}$.
The formal statement of the result in Ref. [114] is summarized as follows.
Lemma E.1 (Adapted from Lemma 1, [114]).
Define the concept class as
$\mathcal{F}_{\mathsf{Hard}}=\left\{f(\bm{x}^{\perp})=\operatorname{Tr}\left(U\ket{\bm{x}^{\perp}}\bra{\bm{x}^{\perp}}U^{\dagger}O\right)\Big{|}\bm{x}^{\perp}\in\{-1,1\}^{N},O\sim\mathbb{D}_{O}\right\},$ (74)
where $\bm{x}^{\perp}$ specifies the initial state |$\ket{\bm{x}^{\perp}}$, $U$ denotes the quantum circuit containing at most a polynomial number of quantum gates and $O$ refers to a linear combination of local Pauli strings sampled from a prior distribution $\mathbb{D}_{O}$.
Under the assumption that there exists a distribution $\mathbb{D}^{\prime}$ such that $\mathsf{BQP}\not\subset\mathsf{HeurP}^{\mathbb{D}^{\prime}}/\mathsf{poly}$, then for any $\mathsf{BQP}$-complete language, there exists an associated input distribution $\mathbb{D}$ such that no randomized algorithm that can solely use the collected measure-out data to output a prediction model $h(\bm{x})$, which learns the concept class $\mathcal{F}_{\mathsf{Hard}}$ in a polynomial time. Note that the learning algorithm is said to successfully learn $\mathcal{F}_{\mathsf{Hard}}$ if for any $f\in\mathcal{F}_{\mathsf{Hard}}$, it can output a prediction model $h(\bm{x})$ that satisfies $\mathbb{E}_{\bm{x}\sim\mathbb{D}}[|f(\bm{x})-h(\bm{x})|^{2}]\leq\epsilon$ with high probability.
It is noteworthy that the terminology ‘learning by measure-out data’ that appeared in the above lemma is identical to the learning paradigm investigated in this work. More precisely, the learner is restricted to collecting information from quantum systems within a polynomial time (i.e., applying a polynomial number of measurements in total). Once the data collection process is completed, the learner utilizes the collected data to infer the prediction model $h(\bm{x})$ on the classical side.
The main idea of the proof for Lemma E.1 is as follows. First, the authors correlate $\mathcal{F}_{\mathsf{Hard}}$ with $\mathsf{BQP}$ circuits. In particular, the unitary in $\mathcal{F}_{\mathsf{Hard}}$ is specified to be a family of quantum circuit $\{U_{\mathsf{BQP}^{N}}\}$, which decides the $\mathsf{BQP}$-complete language $\mathsf{L}$, one circuit per size; the observable is set as $O=Z\otimes I_{2}\otimes\cdots\otimes I_{2}$. In this way, the output of quantum circuits can correctly decide every $\bm{x}^{\perp}\in\mathsf{L}$, i.e., $f(\bm{x}^{\perp})>0$ if $\bm{x}\in\mathsf{L}$ and $f(\bm{x}^{\perp})<0$ if $\bm{x}\notin\mathsf{L}$. Then, the authors show that if there exists a classical algorithm that solely utilizes the measure-out data to learn $\mathcal{F}_{\mathsf{Hard}}$ in a polynomial time, then such algorithm could decide any $\mathsf{BQP}$ language on average, which implies $\mathsf{BQP}\subseteq\mathsf{HeurP}^{\mathbb{D}}/\mathsf{poly}$ for any $\mathbb{D}$. However, under widely believed assumptions, for any $\mathsf{BQP}$ complete language $\mathsf{L}$ there always exists a distribution such that $(\mathsf{L},\mathbb{D})\not\subset\mathsf{HeurP}/\mathsf{poly}$, indicating such classical algorithm cannot exist.
@@
SM F Learnability of the proposed kernel-based ML model (Proof of Theorem II.2)
@@
This section provides the proof of Theorem II.2, which analyzes how the prediction error of the proposed kernel-based ML model depends on the number of training examples $n$, the size of the quantum system $N$, and the dimension of classical inputs $d$. Recall that in the main text, the proposed state prediction model is
@@
$\hat{\sigma}_{n}(\bm{x})=\frac{1}{n}\sum_{i=1}^{n}\kappa_{\Lambda}\left(\bm{x},\bm{x}^{(i)}\right)\tilde{\rho}_{T}(\bm{x}^{(i)})\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \kappa_{\Lambda}\left(\bm{x},\bm{x}^{(i)}\right)=\sum_{\bm{\omega},\|\bm{\omega}\|_{0}\leq\Lambda}2^{\|\bm{\omega}\|_{0}}\Phi_{\bm{\omega}}(\bm{x})\Phi_{\bm{\omega}}(\bm{x}^{(i)})\in\mathbb{R}.$ (75)
@@
What we intend to prove is the average discrepancy between $\operatorname{Tr}(\hat{\sigma}_{n}(\bm{x})O)$ and the ground truth $\operatorname{Tr}(\rho(\bm{x})O)$ when $\bm{x}$ is uniformly and randomly sampled from $[-\pi,\pi]^{d}$ and the local observable $O$ is sampled from a prior distribution $\mathbb{D}_{O}$, i.e., $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}|\operatorname{Tr}(O\hat{\sigma}_{n}(\bm{x}))-\operatorname{Tr}(O\rho(\bm{x}))|^{2}$.
@@
Before moving to proceed with the further analysis, let us exhibit the formal statement of Theorem II.2.
@@
Following notations in the main text, consider a parametrized family of $N$-qubit states $\mathcal{Q}$ and a sum $O=\sum_{i=1}O_{i}$ of multiple local observables with $\sum_{i}\|O_{i}\|_{\infty}\leq B$ and the maximum locality of $\{O_{i}\}$ being $K$. Suppose $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}\leq C$. Then, let the dataset be $\mathcal{T}_{\mathsf{s}}=\{\bm{x}^{(i)}\rightarrow\tilde{\rho}(\bm{x}^{(i)})\}_{i=1}^{n}$ with $\bm{x}^{(i)}\sim\textnormal{Unif}[-\pi,\pi]^{d}$ and $n=|\mathfrak{C}(\Lambda)|\frac{2B^{2}9^{K}}{\epsilon}\log\left({2\cdot|\mathfrak{C}(\Lambda)|}/{\delta}\right)$ with $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. When the frequency is truncated to $\Lambda=4C/\epsilon$, the state prediction model in Eq. (75) achieves
@@
To reach Theorem II.2, we first use the triangle inequality to decouple the difference between the prediction and ground truth into the truncation error and the estimation error, i.e.,
@@
$\leq$ $\Bigg{(}\sqrt{\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left[\left|\operatorname{Tr}\left(O\rho_{\Lambda}(\bm{x})\right)-\operatorname{Tr}(O\rho(\bm{x}))\right|^{2}\right]}+\sqrt{\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left[\left|\operatorname{Tr}\left(O\sigma_{n}(\bm{x})\right)-\operatorname{Tr}(O\rho_{\Lambda}(\bm{x}))\right|^{2}\right]}\Bigg{)}^{2}.$ (78b)
@@
Following notations in Theorem II.2, assuming $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}\leq C$, the truncation error induced by removing high-frequency terms of $\rho$ under the trigonometric expansion with $\|\bm{\omega}\|_{0}\leq\Lambda$ is upper bounded by
@@
Following notations in Theorem II.2, with probability at least $1-\delta$, the estimation error induced by finite training examples $\mathcal{T}=\{\tilde{\rho}(\bm{x}^{(i)})\}_{i=1}^{n}$ is upper bounded by
@@
where $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \|\bm{\omega}\|_{0}\leq\Lambda\}$ refers to the set of truncated frequencies.
We are now ready to present the proof of Theorem II.2.
Proof of Theorem II.2.

### v2

@@ A.4 Literature review
    The training dataset $\mathcal{T}=\{(\bm{x}^{(i)},\hat{\bm{o}}^{(i)}_{\bm{a}^{*}})\}_{i=1}^{n}$ with the size $n$ is collected based on $f_{\bm{a}^{*}}$, i.e., for the $i$-th training example, the classical input $\bm{x}^{(i)}$ randomly sampled from $\pi,\pi]^{d}$ is fed into the quantum system described by $f_{\bm{a}^{*}}$ and the statistical estimation of $f_{\bm{a}^{*}}(\bm{x}^{(i)},O)$ is obtained by $T$ shots with $\hat{\bm{o}}^{(i)}_{\bm{a}^{*}}=\frac{1}{T}\sum_{t}\hat{\bm{o}}^{(i)}_{\bm{a}^{*},t}$ and $\mathbb{E}(\hat{\bm{o}}^{(i)}_{\bm{a}^{*}})=f_{\bm{a}^{*}}(\bm{x}^{(i)},O)$;
@@
Remark. For ease of analysis, in the proof of Theorem B.1, we consider the shot number $T$ is sufficiently large such that the measured results can be approximated by the normal distribution with the mean $\mu_{\bm{a}}=f_{\bm{a}}(\bm{x})$. This assumption is widely utilized in quantum computation and quantum information theory [111, 112, 113, 114]. Besides, for each $\bm{a}$, the corresponding variance of measured results is assumed to be equal with the varied $\bm{x}$, i.e., for all $\bm{x}$, the variance is $\nu_{\bm{a}}=\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\nu_{\bm{a}}(\bm{x})$. Note that the assumption can be omitted if the primary focus is solely on the number of $\operatorname{RZ}$ gates $d$, without considering the measurement cost $T$. In this case, the upper bound of the mutual information $I(\bm{a}^{*};\bar{\bm{a}})$ can be effectively derived using Holevo’s theorem [115, 116], where a similar lower bound of the sample complexity can still be obtained with a linear dependence on $d$.
@@
For the generic variance $\nu^{2}$, it is derived as follows. When the state $\rho_{\bm{a}}(\bm{x})$ is measured by $O=BX$, the probability of measuring the state associated with the eigenvalue +1+1 and the state associated with the eigenvalue 1-1 is
@@
Following the results [117, Lemma 9], the right-hand side term is upper bounded by the number of training examples and the qubit count, i.e.,
@@
Consider a learner quires an $N$-qubit quantum circuits and collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[-\pi,\pi]^{d}$. Suppose with high probability, there exists a learned model $h(\cdot,\cdot)$ that can achieve
@@
$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}(\rho(\bm{x})O)\Big|\bm{x}\in[-\pi,\pi]^{d},\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})\right\},$ (65)
where $\bm{x}\in[-\pi,\pi]^{d}$ is the classical input control, $\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})$ refers to the set of circuit layouts formed by $d$ $\operatorname{RZ}$ gates and GdG-d Clifford gates, and $\rho(\bm{x})$ is the quantum state generated by the bounded gates with $\bm{x}$, i.e., $\rho(\bm{x})=U(\bm{x})\rho_{0}U(\bm{x})^{\dagger}$ and the layout of $U(\bm{x})$ follows an element of $\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})$. The diversity of circuit layouts stems from the constraint that the learner can solely utilize the classical input and the corresponding measurement results to infer the target concept, where prior information regarding the quantum circuit indicates that it comprises $G$ gates, with $d$ $\operatorname{RZ}$ gates and dGd-G CI gates. From the perspective of the learner, the quantum computer can yield many circuit layouts, leading that the function space $\mathcal{F}$ contains in total $|\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})|$ circuit layouts.
@@
$\mathbb{E}_{\bm{x}\sim\mathbb{D}_{\mathcal{X}}}\left|\hat{h}_{\mathcal{T}}(\bm{x})-f^{*}(\bm{x})\right|^{2}\leq 12\epsilon\quad\text{with probability at least $1-\delta$}.$ (70)
Remark. Note that although the original proof of Proposition D.1 only concerns $\|O\|\leq 1$, the pertinent proof can be readily extended to the setting of $\sum_{i}\|O_{i}\|_{\infty}\leq B$ concerned in Theorem 1, and therefore we omit it here. Moreover, it only focuses on the case of $T=1$, but the results still hold for the case of $T>1$. This is because in the extreme case with $T\rightarrow\infty$, the required training data size is reduced to $n\geq{38B^{2}\log(2\mathcal{M}(\mathcal{F},4\epsilon,\mathfrak{d})/\delta)}/{\epsilon}$ [68, Equation (C66)]. In this regard, the setting of $T\geq 1$ only trivially influences the sample complexity bound (at most logarithmically). For this reason, we omit the relevant analysis in our proof.
@@
We now derive the upper bound of $|\mathsf{Arc}(\operatorname{RZ},\operatorname{\text{CI}})|$. The total number of possible layouts for an $N$-qubit circuit consisting with $d$ $\operatorname{RZ}$ gates and GdG-d CI gates (i.e., $\text{CI}=\{H,S,\operatorname{CNOT}\}$) is
@@
where the first term $\binom{G}{d}$ computes the number of different arrangements for placing $\operatorname{RZ}$ or $\operatorname{\text{CI}}$ gates at each circuit depth, the second term $N^{d}$ calculates the total combinations of placing $d$ $\operatorname{RZ}$ gates on different qubit wires, the third term $3^{G-d}$ counts the total combinations of choosing different gates from the $\operatorname{\text{CI}}$ gate set, and the last term $\binom{N}{2}^{G-d}$ calculates the upper bound for the total combinations of placing the selected $\operatorname{\text{CI}}$ gates on different qubit wires.
@@
$\mathsf{BQP}$ complexity class. $\mathsf{BQP}$ (short for bounded-error quantum polynomial time) is the class of promise problems that can be solved by polynomial-time quantum computations that may have some small probability of making an error [122]. Let $A=(A_{\text{yes}},A_{\text{no}})$ be a promise problem and let $a,b:\mathbb{N}\rightarrow[0,1]$ be functions. Then $A\in\mathsf{BQP}(a,b)$ if and only if there exists a polynomial-time generated family of quantum circuits $\mathcal{Q}=\{\mathcal{Q}_{N}:N\in\mathbb{N}\}$, where each circuit $\mathcal{Q}_{N}$ takes $N$ input qubits and produces one output qubit, that satisfies the following properties: (1) if $\bm{x}\in A_{\text{yes}}$ then $\Pr(\mathcal{Q}\penalty 10000\ \text{accepts}\penalty 10000\ \bm{x})\geq a(|\bm{x}|)$, and (2) if $\bm{x}\in A_{\text{no}}$ then $\Pr(\mathcal{Q}\penalty 10000\ \text{accepts}\penalty 10000\ \bm{x})\leq b(|\bm{x}|)$. The class is defined as $\mathsf{BQP}=\mathsf{BQP}(2/3,1/3)$.
$\mathsf{\mathsf{P/poly}}$ complexity class. $\mathsf{\mathsf{P/poly}}$, different from $\mathsf{BQP}$, represents the class of decision problems that can be solved by a polynomial-time deterministic algorithm with access to an auxiliary ‘advice’ bitstring [123]. This advice bitstring depends only on the input size and must be identical for all inputs of a given size, regardless of the specific input $\bm{x}$. More specifically, a language $\mathsf{L}:\{0,1\}^{*}\to\{0,1\}$ is in $\mathsf{P/poly}$ if there exists: (1) a polynomial-time deterministic algorithm $\mathcal{A}$, and (2) a family of ‘advice’ bitstrings $\{\mathsf{a}_{N}\}_{N\in\mathbb{N}}$, where $\mathsf{a}_{N}\in\{0,1\}^{\text{poly}(N)}$, such that for every input $\bm{x}\in\{0,1\}^{N}$: $\mathcal{A}(\bm{x},\mathsf{a}_{N})=\mathsf{L}(\bm{x})$. Here $\{0,1\}^{*}$ denotes the set of all finite-length binary strings.
Note that it is widely believed that $\mathsf{BQP}$ is not entirely contained within $\mathsf{P/poly}$. Otherwise, problems like factoring would become solvable by polynomial-size classical circuits.
Learning procedure. Ref. [72] leveraged the relation between $\mathsf{BQP}$ and $\mathsf{P/poly}$ classes to understand the fundamental limitation of classical learners when learning observables from the measurement data. Denote the concept class as
$\mathcal{F}=\left\{f_{\bm{\alpha}}(\bm{x}^{\perp})=\operatorname{Tr}\left(\rho_{\mathsf{H}}(\bm{x}^{\perp})O(\bm{\alpha})\right)\Big|\bm{\alpha}\in\{-1,1\}^{q}\right\},$ (74)
@@
The goal of the classical learner is to learn a model $h(\bm{x}^{\perp})$ which approximates the unknown concept $f_{\bm{\alpha}}(\bm{x}^{\perp})$. The model $h(\bm{x}^{\perp})$ is inferred by the training dataset $\mathcal{T}=\{\bm{x}^{\perp,(i)},\hat{y}^{(i)}\}_{i=1}^{n}$ where $\hat{y}^{(i)}$ refers to the estimated label of $y^{(i)}=\operatorname{Tr}(\rho_{\mathsf{H}}(\bm{x}^{\perp,(i)})O(\bm{\alpha}))$ with an additive error $\epsilon_{2}$ for $\forall i\in[n]$. The estimation error arises from the finite number of measurements, as the classical learner inputs data $\bm{x}$ into the quantum system and relies on a limited number of measurements to obtain the estimated value. An immediate observation is that the procedure of learning by measure-out data is identical to the learning paradigm investigated in this work. Once the data collection process is completed, the learner utilizes $\mathcal{T}$ to infer the prediction model $h(\bm{x}^{\perp})$ on the classical side.
The efficacy of the classical learner towards the concept class $\mathcal{F}$ is quantified by the following definition.
Definition 2 (Efficient learning condition, Definition 4 in Ref. [72]).
A concept class $\mathcal{F}$ in Eq. (74) is efficiently learnable if there exists a polynomial-time algorithm $\mathcal{A}$ with complexity $\text{poly}(1/\epsilon,1/\delta,1/\epsilon_{2},N)$, such that for any $\epsilon,\epsilon_{2}>0$, any $0<\delta<1$, any target function $f_{\alpha}\in\mathcal{F}$, and any input distribution $\mathbb{D}$, the following holds:
If $\mathcal{A}$ receives a training set $\mathcal{T}=\{\bm{x}^{\perp,(i)},\hat{y}^{(i)}\}_{i=1}^{n}$ with $n\sim\text{poly}(1/\epsilon,1/\delta,1/\epsilon_{2},N)$, it produces a hypothesis $h(.)=\mathcal{A}(\mathcal{T},\epsilon,\delta,.)$ such that, with probability at least $1-\delta$, the expected error satisfies $\mathbb{E}_{\mathbf{\bm{x}^{\perp}}\sim\mathbb{D}}\left[|f_{\bm{\alpha}}(\bm{x}^{\perp})-h(\bm{x}^{\perp})|^{2}\right]\leq\epsilon.$
By leveraging the concept of $\mathsf{BQP}$ and $\mathsf{P/poly}$ classes, Ref. [72] shows the hardness of learning $\mathcal{F}$ from the measurement data.
Lemma E.1 (Adapted from Theorem 2, [72]).
For any $\mathsf{BQP}$-complete language, there exists a Hamiltonian $\mathsf{H}_{\mathsf{hard}}$ to build the concept class
$\mathcal{F}_{\mathsf{Hard}}=\left\{f_{\bm{\alpha}}(\bm{x}^{\perp})=\operatorname{Tr}\left(\rho_{\mathsf{H}_{\mathsf{hard}}}(\bm{x}^{\perp})O(\bm{\alpha})\right)\Big|\bm{\alpha}\in\{-1,1\}^{q}\right\},$ (75)
[…]
The main idea of the proof for Lemma E.1 is as follows. First, the authors correlate $\mathcal{F}_{\mathsf{Hard}}$ with $\mathsf{BQP}$ circuits. In particular, the unitary $U_{\mathsf{H}_{\mathsf{hard}}}$ in $\mathcal{F}_{\mathsf{Hard}}$ is specified to be a family of quantum circuit $\{U_{\mathsf{BQP}^{N}}\}_{N}$, which decides the $\mathsf{BQP}$-complete language $\mathsf{L}$, one circuit per size; the observable is set as $O(\bm{\alpha}^{*})=Z\otimes\mathbb{I}_{2}\otimes\cdots\otimes\mathbb{I}_{2}$. In this way, the output of quantum circuits can correctly decide every $\bm{x}^{\perp}\in\mathsf{L}$, i.e., $f(\bm{x}^{\perp})>0$ if $\bm{x}\in\mathsf{L}$ and $f(\bm{x}^{\perp})<0$ if $\bm{x}\notin\mathsf{L}$. Then, the authors show that if $\mathcal{F}_{\mathsf{Hard}}$ is learnable in the sense of Def. 2, then there must exists a polynomial size classical circuit which evaluates $f_{\bm{\alpha}^{*}}=\operatorname{Tr}(\rho_{\mathsf{H}_{\mathsf{hard}}}(\bm{x}^{\perp})O(\bm{\alpha}^{*}))$ correctly on every $\bm{x}^{\perp}\in\{0,1\}^{N}$, implying $\mathsf{BQP}\subseteq\mathsf{P/poly}$.
We next generalize the results of Lemma E.1 to show the computational hardness of learning the bounded-gate circuit with incoherent dynamics. To achieve this goal, it is sufficient to show the concept class $\mathcal{F}$ in Eq. (3) of the main text can represent the concept class of $\mathsf{BQP}$ circuits discussed above.
[…]
The reformulation enables us to generalize the hardness of learning quantum observables from measured data to the broader challenge of learning bounded-gate circuits, which is the focus of our work. That is, according to the results of Lemma E.1, we can conclude that no algorithm can only use the measure-out data to learn the bounded-gate quantum circuit within a polynomial time unless $\mathsf{BQP}\subseteq\mathsf{P/poly}$. This proves Theorem B.3.
SM F Learnability of the proposed kernel-based ML model (Proof of Theorem 2)
@@
This section provides the proof of Theorem 2, which analyzes how the prediction error of the proposed kernel-based ML model depends on the number of training examples $n$, the size of the quantum system $N$, and the dimension of classical inputs $d$. Recall that in the main text, the proposed state prediction model is
@@
$\hat{\sigma}_{n}(\bm{x})=\frac{1}{n}\sum_{i=1}^{n}\kappa_{\Lambda}\left(\bm{x},\bm{x}^{(i)}\right)\tilde{\rho}_{T}(\bm{x}^{(i)})\penalty 10000\ \text{with}\penalty 10000\ \kappa_{\Lambda}\left(\bm{x},\bm{x}^{(i)}\right)=\sum_{\bm{\omega},\|\bm{\omega}\|_{0}\leq\Lambda}2^{\|\bm{\omega}\|_{0}}\Phi_{\bm{\omega}}(\bm{x})\Phi_{\bm{\omega}}(\bm{x}^{(i)})\in\mathbb{R}.$ (76)
@@
What we intend to prove is the average discrepancy between $\operatorname{Tr}(\hat{\sigma}_{n}(\bm{x})O)$ and the ground truth $\operatorname{Tr}(\rho(\bm{x})O)$ when $\bm{x}$ is uniformly and randomly sampled from $\pi,\pi]^{d}$ and the local observable $O$ is sampled from a prior distribution $\mathbb{D}_{O}$, i.e., $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}|\operatorname{Tr}(O\hat{\sigma}_{n}(\bm{x}))-\operatorname{Tr}(O\rho(\bm{x}))|^{2}$.
@@
Before moving to proceed with the further analysis, let us exhibit the formal statement of Theorem 2.
@@
Following notations in the main text, consider a parametrized family of $N$-qubit states $\mathcal{Q}$ and a sum $O=\sum_{i=1}O_{i}$ of multiple local observables with $\sum_{i}\|O_{i}\|_{\infty}\leq B$ and the maximum locality of $\{O_{i}\}$ being $K$. Suppose $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}\leq C$. Then, let the dataset be $\mathcal{T}_{\mathsf{s}}=\{\bm{x}^{(i)}\rightarrow\tilde{\rho}(\bm{x}^{(i)})\}_{i=1}^{n}$ with $\bm{x}^{(i)}\sim\textnormal{Unif}[-\pi,\pi]^{d}$ and $n=|\mathfrak{C}(\Lambda)|\frac{2B^{2}9^{K}}{\epsilon}\log\left({2\cdot|\mathfrak{C}(\Lambda)|}/{\delta}\right)$ with $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\penalty 10000\ s.t.\penalty 10000\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. When the frequency is truncated to $\Lambda=4C/\epsilon$, the state prediction model in Eq. (76) achieves
@@
To reach Theorem 2, we first use the triangle inequality to decouple the difference between the prediction and ground truth into the truncation error and the estimation error, i.e.,
@@
$\leq$ $\Bigg(\sqrt{\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left[\left|\operatorname{Tr}\left(O\rho_{\Lambda}(\bm{x})\right)-\operatorname{Tr}(O\rho(\bm{x}))\right|^{2}\right]}+\sqrt{\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left[\left|\operatorname{Tr}\left(O\hat{\sigma}_{n}(\bm{x})\right)-\operatorname{Tr}(O\rho_{\Lambda}(\bm{x}))\right|^{2}\right]}\Bigg)^{2}.$ (79b)
@@
Following notations in Theorem 2, assuming $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}\leq C$, the truncation error induced by removing high-frequency terms of $\rho$ under the trigonometric expansion with $\|\bm{\omega}\|_{0}\leq\Lambda$ is upper bounded by
@@
Following notations in Theorem 2, with probability at least $1-\delta$, the estimation error induced by finite training examples $\mathcal{T}=\{\tilde{\rho}(\bm{x}^{(i)})\}_{i=1}^{n}$ is upper bounded by
@@
where $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\penalty 10000\ s.t.\penalty 10000\ \|\bm{\omega}\|_{0}\leq\Lambda\}$ refers to the set of truncated frequencies.
We are now ready to present the proof of Theorem 2.
Proof of Theorem 2.